# Multiplier ideals of normal surface singularities

### v1

@@ Abstract.
We study the multiplier ideals and the corresponding jumping numbers and multiplicities $\{m(c)\}_{c\in\mathbb{R}}$ in the following context: $(X,o)$ is a complex analytic normal surface singularity, ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$ is an ${\mathfrak{m}}_{X,o}$–primary ideal, $\phi:\widetilde{X}\to X$ is a log resolution of $\mathfrak{a}$ such that $\mathfrak{a}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$, for some nonzero effective divisor $F$ supported on $\phi^{-1}(0)$.
We show that $\{m(c)\}_{c>0}$ is combinatorially computable from $F$ and the resolution graph $\Gamma$ of $\phi$, and we provide several formulae. We also extend Budur’s result (valid for $(X,o)=(\mathbb{C}^{2},0)$), which makes an identification of $\sum_{c\in[0,1]}m(c)t^{c}$ with a certain Hodge spectrum. In our general case we use Hodge spectrum with coefficients in a mixed Hodge module.
We show that $\{m(c)\}_{c\leq 0}$ usually depends on the analytic type of $(X,o)$. However, for some distinguished analytic types we determine it concretely. E.g., when $(X,o)$ is weighted homogeneous (and $F$ is associated with the central vertex), we recover $\sum_{c}m(c)t^{c}$ from the Poincaré series of $(X,o)$ and when $(X,o)$ is a splice quotient then we recover $\sum_{c}m(c)t^{c}$ from the multivariable topological Poincaré (zeta) function of $\Gamma$.
@@ 1. Introduction
Assume that $(X,o)$ is a complex analytic normal surface singularity, let ${\mathscr{O}}_{X,o}$ be its local algebra with maximal ideal ${\mathfrak{m}}_{X,o}$. Let us also fix an ${\mathfrak{m}}_{X,o}$–primary ideal ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$. Then there exists a log resolution $\phi:\widetilde{X}\to X$ and a nonzero integral cycle $F$ supported on the exceptional curve $\phi^{-1}(0)=E$ such that ${\mathfrak{a}}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$. In fact, in our study we will start with some $\phi$ and such a cycle $F$.
We denote by $Z_{K}$ the anticanonical cycle associated with $\phi$, $-c_{1}(\Omega^{2}_{\widetilde{X}})=-c_{1}({\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}}))\in H^{2}(\widetilde{X},\mathbb{Z})$, identified with a rational cycle (via the adjunction relations). We consider for any $c\in\mathbb{R}$ the cycle $L(c):=\lfloor Z_{K}+cF\rfloor$. Then the tower of multiplier ideals is defined as $\mathfrak{I}(c,F):=\phi_{*}{\mathscr{O}}_{\widetilde{X}}(-L(c))\subset{\mathscr{O}}_{X,o}$. They measure the complexity of the singularity $(X,o)$ compatibly with $F$ (or with $\mathfrak{a}$). They were introduced in [Ein, Laz2] and since then they were intensively studied, see [budur, Carrami, Carrami2, Galindo, Galindo2, Howald, Hyry, Jarvilehto, Montaner, Naie, Tucker] with all their connections with different aspects of singularities and singular map-germs. In fact, the definition and study extends to any dimension, see e.g. [Ein, Laz2, Pande] and the references therein.
It turns out that for any $c\in\mathbb{R}$, the ideal $\mathfrak{I}(c,F)\subset{\mathscr{O}}_{X,o}$ has finite codimension, $\mathfrak{I}(c,F)\subset\mathfrak{I}(c^{\prime},F)$ for $c^{\prime}<c$, and $\mathfrak{I}(c,F)={\mathscr{O}}_{X,o}$ for $c\ll 0$. The number $c$ is called jumping number if $m(c):=\dim\,\mathfrak{I}(c-\epsilon,F)/\mathfrak{I}(c,F)\not=0$ for $0<\epsilon\ll 1$, and $m(c)$ is the corresponding jumping multiplicity. One verifies that in any compact interval there are only finitely many jumping numbers and they are all rational.
@@
(B)  In the earlier studies only the cases $c>0$ were treated. Our study of the cases $c\leq 0$ opens a new perspectives. If $(X,o)$ is smooth or log terminal then $m(c)=0$ for $c\leq 0$, however, in the general case (even for rational $(X,o)$) the negative zone appears as a very challenging part.
In fact (and this is one of the main messages of the note), we show that the study of the Poincaré series of multiplier ideals $\sum_{c\in\mathbb{R}}m(c)t^{c}$ (introduced first in [Galindo]) breaks into two major parts with rather different aspects.
@@
Regarding this we stress out two facts. Firstly, Budur in [budur] identified $\sum_{c\in(0,1]}m(c)t^{c}$ with the Hodge spectrum ${\rm Sp}_{(0,1]}(f)$ of a function germ $f$, where in his case, $(X,o)=(\mathbb{C}^{2},0)$ is smooth, $f:(\mathbb{C}^{2},0)\to(\mathbb{C},0)$ is a function germ with an isolated singularity, $\phi$ is a log embedded resolution and $F$ is the divisor ${\rm div}_{\phi,E}(f)$ of $\phi^{*}(f)$ along $\phi^{-1}(0)=E$. This was reproved in [Carrami] as well. (For the theory of mixed Hodge structures and of the spectrum associated with isolated hypersurface singularities see the work of Steenbrink, e.g. [steenbrink, steenbrink2] or [SSS].)
@@
This formula will resonate perfectly with Hodge spectrum computations of a generic element $f\in\mathfrak{a}$. Indeed, if $(X,o)$ is numerically Gorenstein (that is, $Z_{K}$ is an integral cycle), and $f$ is an analytic function gems $(X,o)\to(\mathbb{C},0)$ with isolated singularities, ${\rm div}_{\phi,E}(f)=F$, we prove the generalization of the statement of Budur: $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(t)$, cf. Corollary LABEL:cor:NS. However, when $\Gamma$ is not numerically Gorenstein, $\{Z_{K}\}\not=0$, then we need to use a more general spectrum computation established in [NS]. In [NS] the usual spectrum is replaced by the spectrum of $f$ with coefficients in a mixed Hodge module ${\mathscr{M}}$ of $(X,o)$, such that ${\mathscr{M}}$ restricted to $X\setminus\{o\}$ is a variation of Hodge structure whose underlying representation is abelian. Indeed, under the assumption that the link is a rational homology sphere, from $Z_{K}$ we construct a character of the first homology of the link, hence a rank one variation of Hodge structure ${\mathbb{V}}$ on $X\setminus\{o\}$, and we prove the identity $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(t)(f,j_{*}{\mathbb{V}})$ (where $j:X\setminus\{o\}\subset X$ is the inclusion), see Theorem LABEL:th:NS2. The construction of ${\mathbb{V}}$ can be done using the canonical cyclic covering of $(X,o)$ (associated with $Z_{K}$, or in $\mathbb{Q}$–Gorenstein case associated with the canonical divisor $K_{\widetilde{X}}$, as well). We emphasize again, though a priori both objects $J_{[0,1]}(t)$ and ${\rm Sp}_{[0,1]}(t)(f,j_{*}{\mathbb{V}})$ are analytical, we provide combinatorial formulae for them.
@@
The second new aspect regarding $\sum_{c\in(0,1]}m(c)t^{c}$ is the following connection. Assume that there exists again a function $f:(X,o)\to(\mathbb{C},0)$ with an isolated singularity such that $F={\rm div}_{\phi,E}(f)$. Set $(C,o):=(f^{-1}(0),o)$ and let $\delta(C,o)$ be its delta invariant. Then $\sum_{c\in(0,1]}m(c)=\delta(C,o)+(\{Z_{K}\},F)$.
@@
(ii)  (the $c\leq 0$ zone)  We show by examples that the multiplicity $m(0)$ cannot be determined topologically. However, we characterize it completely in terms of the analytic type of $(X,o)$. The analytic contribution is a term $\mathfrak{h}^{0}\in\{0,1\}$ characterized by an $h^{0}$–sheaf cohomology.
@@
The $c\leq 0$ zone can be called the ‘$h^{0}$–zone’, where the Riemann–Roch expressions or vanishing theorems cannot be applied. In this zone $m(c)$ usually cannot be computed form $\Gamma$ and $F$. We exemplify this fact by concrete examples when we fix $\Gamma$ and $F$ and we consider several different analytic realizations, and we compute the $m(c)$’s. Interestingly enough, already the numerical invariant $\sum_{c\leq 0}m(c)$ measures the analytic type and it is connected to the most important analytic invariant of $(X,o)$, the geometric genus $p_{g}(X,o)$. More precisely, we prove that $\sum_{c\leq 0}m(c)=p_{g}(X,o)+\chi(\lfloor Z_{K}\rfloor)$. Already when $(X,o)$ is rational (i.e. $p_{g}(X,o)=0$), $\sum_{c\leq 0}m(c)$ can be arbitrarily large (see e.g. Example LABEL:ex:sum).
@@
$\bullet$  If the link is a rational homology sphere and we fix a resolution graph $\Gamma$, and we assume that the analytic type of $(X,o)$ is splice quotient with respect to $\Gamma$ (a family introduced by Neumann and Wahl, cf. [NW]). Then in Proposition LABEL:PZ we provide a combinatorial formula for $\sum_{c\in\mathbb{R}}m(c)t^{c}$ in terms of $\{L(c)\}_{c}$ and the multivariable topological Poincaré (zeta) function $Z({\bf t})$ of $\Gamma$. This includes e.g. all the cases when $(X,o)$ is rational or minimally elliptic (with $\Gamma$ minimal). (For more on $Z({\bf t})$ see [CDG, CDGEq, Nkonyv].)
But we emphasize again that in general, the computation of $\{m(c)\}_{c\leq 0}$ is of analytic nature, hence it runs in parallel with the classification of the analytic types of $(X,o)$ supported of a fixed topological type identified by $\Gamma$. We will return back to this study in the forthcoming manuscript.
@@ 2.1. Notations regarding resolutions and resolution graphs
We denote by $M$ the link of $(X,o)$. It is known that $M$ is a rational homology sphere if and only if each $E_{i}$ is rational and $\Gamma$ is a tree. It is an integral homology sphere if additionally $\det\,(\,,\,)=\pm 1$. More generally, if $g_{i}$ denotes the genus of $E_{i}$, $g:=\sum_{i}g_{i}$ and $h$ the first Betti number of the topological realization of $\Gamma$ (the number of independent cycle in the graph), then the first Betti number of $M$ is $2g+h$, and ${\rm Tors}\,H_{1}(M,\mathbb{Z})\simeq L^{\prime}/L$.

### v2

@@ Abstract.
We study the multiplier ideals and the corresponding jumping numbers and multiplicities $\{m(c)\}_{c\in\mathbb{R}}$ in the following context: $(X,o)$ is a complex analytic normal surface singularity, ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$ is an ${\mathfrak{m}}_{X,o}$–primary ideal, $\phi:\widetilde{X}\to X$ is a log resolution of $\mathfrak{a}$ such that $\mathfrak{a}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$, for some nonzero effective divisor $F$ supported on $\phi^{-1}(0)$. A priori, $\{m(c)\}_{c\in\mathbb{R}}$ depends on the Hilbert function associated with the resolution and $F$.
We prove that $\{m(c)\}_{c>0}$ is combinatorially computable from $F$ and the resolution graph $\Gamma$ of $\phi$, and we provide several formulae. We also extend Budur’s result (valid for $(X,o)=(\mathbb{C}^{2},0)$), which makes an identification of $\sum_{c\in[0,1]}m(c)t^{c}$ with a certain Hodge spectrum. In our general case we use Hodge spectrum with coefficients in a mixed Hodge module.
We show that $\{m(c)\}_{c\leq 0}$ usually depends on the analytic type of $(X,o)$. However, for some distinguished analytic types we determine it concretely. E.g., when $(X,o)$ is weighted homogeneous (and $F$ is associated with the central vertex), we recover $\sum_{c}m(c)t^{c}$ from the Poincaré series of $(X,o)$. Also, when $(X,o)$ is a splice quotient then we recover $\sum_{c}m(c)t^{c}$ from the multivariable topological Poincaré (zeta) function of $\Gamma$.
@@ 1. Introduction
Assume that $(X,o)$ is a complex analytic normal surface singularity, let ${\mathscr{O}}_{X,o}$ be its local algebra with maximal ideal ${\mathfrak{m}}_{X,o}$. Let us also fix an ${\mathfrak{m}}_{X,o}$–primary ideal ${\mathfrak{a}}\subset{\mathscr{O}}_{X,o}$. Then there exists a log resolution $\phi:\widetilde{X}\to X$ and a nonzero integral cycle $F$ supported on the exceptional curve $\phi^{-1}(0)=E$ such that ${\mathfrak{a}}{\mathscr{O}}_{\widetilde{X}}={\mathscr{O}}_{\widetilde{X}}(-F)$. In fact, in our study we will start with some log resolution $\phi$ and such a nonzero antinef cycle $F$.
We denote by $Z_{K}$ the anticanonical cycle associated with $\phi$, namely $-c_{1}(\Omega^{2}_{\widetilde{X}})=-c_{1}({\mathscr{O}}_{\widetilde{X}}(K_{\widetilde{X}}))\in H^{2}(\widetilde{X},\mathbb{Z})$, identified with a rational cycle (via the adjunction relations). We consider for any $c\in\mathbb{R}$ the cycle $L(c):=\lfloor Z_{K}+cF\rfloor$. Then the tower of multiplier ideals is defined as $\mathfrak{I}(c,F):=\phi_{*}{\mathscr{O}}_{\widetilde{X}}(-L(c))\subset{\mathscr{O}}_{X,o}$. They measure the complexity of the singularity $(X,o)$ compatibly with $F$ (or with $\mathfrak{a}$). They were introduced (at least implicitly) in [[21], [25], [33], [34]] and in the present form in [Ein, Laz2]. Since then they were intensively studied, see [budur, Carrami, Carrami2, Galindo, Galindo2, Howald, Hyry, Jarvilehto, Montaner, Naie, Tucker] with all their connections with different aspects of singularities and singular map-germs. In fact, the definition and study extends to any dimension, see e.g. [Ein, Laz2, Pande] and the references therein.
It turns out that for any $c\in\mathbb{R}$, the ideal $\mathfrak{I}(c,F)\subset{\mathscr{O}}_{X,o}$ has finite codimension, $\mathfrak{I}(c,F)\subset\mathfrak{I}(c^{\prime},F)$ for $c^{\prime}<c$, and $\mathfrak{I}(c,F)={\mathscr{O}}_{X,o}$ for $c\ll 0$. The number $c$ is called jumping number with jumping multiplicity $m(c)$ if $m(c):=\dim\,\mathfrak{I}(c-\epsilon,F)/\mathfrak{I}(c,F)\not=0$ for $0<\epsilon\ll 1$. One verifies that in any compact interval there are only finitely many jumping numbers and they are all rational.
@@
(B)  In the earlier studies only the cases $c>0$ were treated. Our study of the cases $c\leq 0$ opens new perspectives. If $(X,o)$ is smooth or log terminal then $m(c)=0$ for $c\leq 0$, however, in the general case (even for rational $(X,o)$) the negative zone appears as a very challenging part.
In fact (and this is one of the main messages of the note), we show that the study of $\{m(c)\}_{c\in\mathbb{R}}$ (equivalently, the Poincaré series of multiplier ideals $\sum_{c\in\mathbb{R}}m(c)t^{c}$ introduced first in [Galindo]) breaks into two major parts with rather different aspects.
@@
Regarding this we stress out two facts. Firstly, Budur in [budur] identified $\sum_{c\in(0,1]}m(c)t^{c}$ with the Hodge spectrum ${\rm Sp}_{(0,1]}(f,t)$ of a function germ $f$, where in his case, $(X,o)=(\mathbb{C}^{2},0)$ is smooth, $f:(\mathbb{C}^{2},0)\to(\mathbb{C},0)$ is a function germ with an isolated singularity, $\phi$ is a log embedded resolution and $F$ is the divisor ${\rm div}_{\phi,E}(f)$ of $\phi^{*}(f)$ along $\phi^{-1}(0)=E$. This was reproved in [Carrami] as well. (For the theory of mixed Hodge structures and of the spectrum associated with isolated hypersurface singularities see the work of Steenbrink, e.g. [steenbrink, steenbrink2] or [SSS].)
@@
This formula will resonate perfectly with the Hodge spectrum computations of a generic element $f\in\mathfrak{a}$. Indeed, if $(X,o)$ is numerically Gorenstein (that is, $Z_{K}$ is an integral cycle), and $f$ is an analytic function gems $(X,o)\to(\mathbb{C},0)$ with isolated singularities and ${\rm div}_{\phi,E}(f)=F$, we prove the generalization of the statement of Budur: $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(f,t)$, cf. Corollary LABEL:cor:NS. However, when $\Gamma$ is not numerically Gorenstein, $\{Z_{K}\}\not=0$, then we need to use a more general spectrum formula established in [NS]. In [NS] the ‘usual’ spectrum is replaced by the spectrum of $f$ with coefficients in a mixed Hodge module ${\mathscr{M}}$ of $(X,o)$, such that ${\mathscr{M}}$ restricted to $X\setminus\{o\}$ is a variation of Hodge structure whose underlying representation is abelian. Indeed, under the assumption that the link is a rational homology sphere, from $Z_{K}$ we construct a character of the first homology of the link, hence a rank one variation of Hodge structure ${\mathbb{V}}$ on $X\setminus\{o\}$, and we prove the identity $J_{[0,1]}(t)={\rm Sp}_{[0,1]}(f,j_{*}{\mathbb{V}},t)$ (where $j:X\setminus\{o\}\subset X$ is the inclusion), see Theorem LABEL:th:NS2. The construction of ${\mathbb{V}}$ can be done using the canonical cyclic covering of $(X,o)$ (associated with $Z_{K}$, or in $\mathbb{Q}$–Gorenstein case associated with the canonical divisor $K_{\widetilde{X}}$, as well). We emphasize again, though a priori both objects $J_{[0,1]}(t)$ and ${\rm Sp}_{[0,1]}(f,j_{*}{\mathbb{V}},t)$ are analytical, we provide combinatorial formulae for them. It would be interesting to compare these results with the general (Hodge) theoretical construction of [schnell] valid for complex manifolds, and with the Bernstein-Sato polynomial connection from [ELSV].
@@
The second new aspect regarding $\sum_{c\in(0,1]}m(c)t^{c}$ is the following connection. Consider again a function $f:(X,o)\to(\mathbb{C},0)$ with an isolated singularity such that $F={\rm div}_{\phi,E}(f)$. Set $(C,o):=(f^{-1}(0),o)$ and let $\delta(C,o)$ be its delta invariant. Then $\sum_{c\in(0,1]}m(c)=\delta(C,o)+(\{Z_{K}\},F)$.
@@
(ii)  (the $c\leq 0$ zone)  We show by examples that the multiplicity $m(0)$ cannot be determined topologically. However, we characterize it completely in terms of the analytic type of $(X,o)$. The analytic contribution is a term $\mathfrak{h}^{0}\in\{0,1\}$ characterized by an $h^{0}$–sheaf cohomology (see LABEL:s:cpoz).
@@
The $c\leq 0$ zone can be called the ‘$h^{0}$–zone’, where the $h^{1}$–vanishing theorems (hence the Riemann–Roch expressions) cannot be applied. In this zone $m(c)$ usually cannot be computed form $\Gamma$ and $F$. We exemplify this fact by concrete examples when we fix $\Gamma$ and $F$ and we consider several different analytic realizations, and we compute the $m(c)$’s. Interestingly enough, already the numerical invariant $\sum_{c\leq 0}m(c)$ measures the analytic type and it is connected to the most important analytic invariant of $(X,o)$, the geometric genus $p_{g}(X,o)$. More precisely, we prove that $\sum_{c\leq 0}m(c)=p_{g}(X,o)+\chi(\lfloor Z_{K}\rfloor)$. Already when $(X,o)$ is rational (i.e. $p_{g}(X,o)=0$), $\sum_{c\leq 0}m(c)$ can be arbitrarily large (see e.g. Example LABEL:ex:sum).
@@
$\bullet$  In the next case we assume that the link is a rational homology sphere, and we fix a resolution graph $\Gamma$, and we assume that the analytic type of $(X,o)$ is splice quotient with respect to $\Gamma$ (a family introduced by Neumann and Wahl, cf. [NW]). Then in Proposition LABEL:PZ we provide a combinatorial formula for $\sum_{c\in\mathbb{R}}m(c)t^{c}$ in terms of $\{L(c)\}_{c}$ and the multivariable topological Poincaré (zeta) function $Z({\bf t})$ of $\Gamma$. This includes e.g. all the cases when $(X,o)$ is rational or minimally elliptic (with $\Gamma$ minimal). (For more on $Z({\bf t})$ see [CDG, CDGEq, Nkonyv].)
But we emphasize again that in general, the computation of $\{m(c)\}_{c\leq 0}$ is of analytic nature, hence it runs in parallel with the classification of the analytic types of $(X,o)$ supported of a fixed topological type identified by $\Gamma$. We will return back to the study of these analytic aspects in the forthcoming manuscript.
@@ 2.1. Notations regarding resolutions and resolution graphs
We denote by $M$ the link of $(X,o)$. It is known that $M$ is a rational homology sphere if and only if each $E_{i}$ is rational and $\Gamma$ is a tree. It is an integral homology sphere if additionally $\det\,(\,,\,)=\pm 1$. More generally, if $g_{i}$ denotes the genus of $E_{i}$, $g:=\sum_{i}g_{i}$, and $h$ is the first Betti number of the topological realization of $\Gamma$ (the number of independent cycle in the graph), then the first Betti number of $M$ is $2g+h$, and ${\rm Tors}\,H_{1}(M,\mathbb{Z})\simeq L^{\prime}/L$.